OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

OpenStudy (anonymous):

t=4x-2 y=2x-4 y=-3(x-4) y=3(x+4)

OpenStudy (gorv):

y-0=m(x-4)

OpenStudy (gorv):

m=(6-0)/(2-4)=6/-2=-3

OpenStudy (gorv):

y=-3(x-4)

OpenStudy (anonymous):

ahh i see what you did

OpenStudy (anonymous):

i see where i messed up now thanks man!!